OpenStudy (anonymous):

(√8x^2)(√2x^5) A) 2x^3 √2x B) 2x^2 √x C) 3x^4 √x D) 4x^3 √x

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sqrt{8x^2} \cdot \sqrt{2x^5} }\) like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

they confused me alot

OpenStudy (solomonzelman):

You know that; \(\color{#000000 }{ \displaystyle \sqrt{~\rm B~} \cdot \sqrt{~\rm C~}=\sqrt{~\rm B~\cdot~C~}}\)

OpenStudy (solomonzelman):

So, you can use that at first, \(\color{#000000 }{ \displaystyle \sqrt{8x^2} \cdot \sqrt{2x^5}\Longrightarrow \sqrt{16x^7}}\)

OpenStudy (solomonzelman):

are you with me so far?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Ok, good.

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sqrt{16x^7}=\sqrt{16}\cdot \sqrt{x^6}\times \sqrt{x^\color{white}{1}}}\)

OpenStudy (solomonzelman):

can you tell me what the first two terms in the product will simplify to?

OpenStudy (anonymous):

16 and x^6?

OpenStudy (solomonzelman):

what is the square root of 16?

OpenStudy (anonymous):

4

OpenStudy (solomonzelman):

So, so far we can write \(\color{#000000 }{ \displaystyle \sqrt{16x^7}=4\cdot \sqrt{x^6}\times \sqrt{x^\color{white}{1}}}\)

OpenStudy (solomonzelman):

And then, tell me what is \(\color{#000000 }{ \displaystyle \sqrt{x^6}}\) ?? knowing that \(\color{#000000 }{ \displaystyle x^6=x^{3+3}=x^3\cdot x^3=\left(x^3\right)^2}\)

OpenStudy (solomonzelman):

(You know that \(\sqrt{49}=7\), because \(7^2=49\), right?)

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

and so, you can tell me what is \(\color{#000000 }{ \displaystyle \sqrt{x^6}}\), once you know that \(\left(x^3\right)^2=x^6\)

OpenStudy (anonymous):

power of power right

OpenStudy (solomonzelman):

well, usually, (for positive z), you have: \(\color{#000000 }{ \displaystyle \sqrt{z^2}=\sqrt[2]{z^2} }\) (the sides are the same... i.e., when the number in the corner of the root is unspecificed, it is always a 2) \(\color{#000000 }{ \displaystyle \sqrt[2]{z^2} =z^{2/2}}\) (Because there is a rule: \(\color{#000000 }{ \displaystyle \sqrt[n]{a^m} =a^{m/n}}\)) \(\color{#000000 }{ \displaystyle z^{2/2}=z^1=z}\) (because the 2's cancel)

OpenStudy (solomonzelman):

You have the same thing in \(\color{#000000 }{ \displaystyle \sqrt{x^6} }\), except that \(z\), in your case, and as you can easily notice, is \(x^3\).

OpenStudy (solomonzelman):

(I will explain later, why \(x\) is assumed to be positive, in your question)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sqrt{8x^2} \cdot \sqrt{2x^5}= \sqrt{8x^2\cdot 2x^5}}\) \(\color{#000000 }{ \displaystyle \sqrt{8x^2\cdot 2x^5}=\sqrt{16x^7}}\) \(\color{#000000 }{ \displaystyle \sqrt{16x^7}=\sqrt{16}\cdot \sqrt{x^6}\cdot \sqrt{x~}}\) \(\color{#000000 }{ \displaystyle \sqrt{16}\cdot \sqrt{x^6}\cdot \sqrt{x~}=4\cdot \sqrt{x^6}\cdot \sqrt{~x}}\) So so, far we have simplified, till \(\color{#000000 }{ \displaystyle 4\cdot \sqrt{x^6}\cdot \sqrt{~x}}\)

OpenStudy (anonymous):

yes and sry i get confused even if im in 10th with this

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \sqrt{x^6}=\sqrt[2]{x^6}}\) (Remember what I did, \(\sqrt[n]{a^m}=a^{m/n}\)) \(\color{#000000 }{ \displaystyle \sqrt[2]{x^6}=x^{6/2}=x^?}\)

OpenStudy (solomonzelman):

So, tell me \(\color{#000000 }{ \displaystyle \sqrt{x^6}=x^?}\)

OpenStudy (anonymous):

x^3

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

therefore, \(\color{#000000 }{ \displaystyle 4\cdot \sqrt{x^6}\cdot \sqrt{~x}=4\cdot x^3\cdot \sqrt{x~}}\)

OpenStudy (solomonzelman):

So, do you understand all of the steps that we preformed to get the answer?

OpenStudy (anonymous):

i getting it

OpenStudy (anonymous):

it d

OpenStudy (solomonzelman):

Yes

OpenStudy (solomonzelman):

And now I will explain what's not precise about the question

OpenStudy (solomonzelman):

have you heard about the "absolute value" ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Yes, and the absolute value (of z) is defined as, \(\color{#000000 }{ \displaystyle \left|z\right|=\sqrt{z^2}}\) And therefore, \(\color{#000000 }{ \displaystyle \sqrt{x^6}=\sqrt{\left(x^3\right)^2}=\left|x^3\right| }\)

OpenStudy (solomonzelman):

So, really the precise answer is, \(\color{#000000 }{ \displaystyle 4\left|x^3\right|\sqrt{~x}}\)

OpenStudy (solomonzelman):

But, the assumtion made is that \(x\) is NOT NEGATIVE. And therefore, for non-negative \(x\), we have \(|x|=x\) and \(\left|x^3\right|=x^3\).

OpenStudy (anonymous):

can we be friend

OpenStudy (solomonzelman):

What exactly do you mean by being friends? ((I'm usually friendly with people on the site, but actually making "friends" is something I very rarely do))

OpenStudy (anonymous):

me too be i alway like to be friend wit ppl smart than me or nice

OpenStudy (anonymous):

so it a no right

OpenStudy (solomonzelman):

Well, I just hardly imagine online friendships...

OpenStudy (solomonzelman):

usually my friendships are all face to face.

OpenStudy (anonymous):

i did it that how i met my boyfriend thought online and all my friend i had stop being my friend so onlyi hhave online friend and some pal friend

OpenStudy (solomonzelman):

idk ... will see how everything goes.

OpenStudy (anonymous):

ok do u have fb mine candyquinnred

OpenStudy (anonymous):

im a girl also because my picis of me and my bf